# Pressure-temperature phase diagram calculations using polynomial machine learning potentials: A comprehensive study based on global structure prediction and self-consistent phonon calculations

### v1

@@ III.1 Methodology and search space
We employ an RSS method to enumerate globally stable and metastable structures, which corresponds to the multi-start method in global optimization [68]. In this approach, local geometry optimizations are repeatedly performed on random initial structures, and the local minimum structure with the lowest enthalpy is considered the global minimum. This heuristic approach has been previously utilized in the ab initio random structure search (AIRSS) [33, 4].
@@ III.2 Iterative update of MLP
We utilize the optimal MLP described in Sec. II.5 for the global structure search. This MLP is selected from among Pareto-optimal MLPs to balance prediction accuracy for energy and computational efficiency. Although the MLP exhibits high predictive power for various structures, it occasionally predicts incorrect local minimum structures and fails to accurately estimate their energy values.
To enhance the robustness of the global structure search, the MLP is iteratively updated through repeated global structure searches, as follows. Single-point DFT calculations are performed on all local minimum structures identified by the RSS with the MLP, and these results are added to the training dataset. We then retrain the MLP with this expanded dataset and repeat the global structure search using the updated MLP. This iterative approach, which combines the polynomial MLP and RSS, has been previously employed in global structure searches utilizing the polynomial MLP [37, 38, 39].
@@ III.4 Accuracy of updated MLP
Figure 8: Distributions of $E+\bar{\sigma}V$ for local minimum structures, computed using both the MLPs and DFT calculations, at the first and final iterations of the iterative MLP update procedure.
@@
Figure 8 shows the distributions of $E+\bar{\sigma}V$ for local minimum structures obtained in both the first and final iterations, evaluated by the MLPs and DFT calculations. The distribution for each iteration is obtained using the MLP applied at the corresponding iteration. Here, $E$ represents the cohesive energy, $V$ denotes the cell volume, and $\bar{\sigma}$ is the mean normal stress, defined as $\bar{\sigma}=(\sigma_{11}+\sigma_{22}+\sigma_{33})/3$ [69], where $\sigma_{ij}$ denotes the stress component acting in the $j$ direction on a plane normal to the $i$ direction. In the first iteration, many local minimum structures exhibit large prediction errors. However, by incorporating these structures and their DFT-computed properties into the training dataset and retraining the MLP, the predictive accuracy for local minimum structures improves significantly in the final iteration.
@@
Note that the enthalpy of a local minimum structure is given by $E+pV$ when the stress tensor satisfies $\sigma_{ij}=p\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta. In Fig. 8, the stress tensor of the local minimum structure optimized by the MLP is evaluated through a single-point DFT calculation. Because of prediction errors in the MLP, DFT stress tensor components deviate from the condition $\sigma_{ij}=p\delta_{ij}$. Consequently, the DFT enthalpy value of the local minimum structure cannot be directly obtained from a single-point DFT calculation. Therefore, Fig. 8 instead presents the distributions of the alternative measure $E+\bar{\sigma}V$, which serves as an approximate evaluation of the predictive accuracy for enthalpy.
III.5 Globally stable and metastable structures
@@
By performing the RSS with the updated MLP for elemental Si, we enumerate a total of 28,372 local minimum structures. Figure 9(a) presents their relative enthalpy values computed by the MLP. At each pressure, the relative enthalpy is defined with respect to the lowest enthalpy value among the local minimum structures found at the corresponding pressure. The MLP enables the enumeration of numerous structures across all pressure settings. However, the enthalpy values predicted by the MLP still exhibit small but non-negligible prediction errors. Therefore, we compute their enthalpy values using DFT calculations to evaluate the phase stability among these local minimum structures, as follows. First, we select 344 local minimum structures with MLP-calculated relative enthalpies below 30 meV/atom, as shown in Fig. 9(a). We then perform DFT-based geometry optimizations for these structures and eliminate duplicates.
@@
Figure 10: Relative enthalpy values computed using DFT calculations for local minimum structures. Local minimum structures with relative enthalpy values exceeding 10 meV/atom across the entire pressure range are excluded from the analysis. The legend indicates the space groups of the structures or their corresponding prototype structures, and $Z$ denotes the number of atoms in the unit cell. (a) Enthalpy values are shown relative to the Si-V structure for pressures from 0 to 80 GPa in the left panel. The relative enthalpies in the central panel are displayed for the 9–12 GPa range, where multiple structures have very similar enthalpy values. (b) Enthalpy values are shown relative to the $\alpha$-La-type structure for pressures from 80 to 100 GPa.
Table 2 lists the metastable local minimum structures, which correspond to experimentally reported structures [76, 77, 78, 79]. Although the Si-XI structure is not identified in the current global structure search, as mentioned above, all other experimentally reported structures are discovered. We evaluate their enthalpy values using DFT calculations if the MLP predicts enthalpies below 30 meV/atom; otherwise, we use the MLP-calculated values. These structures exhibit positive relative enthalpy values, which is consistent with the fact that they have not been observed as stable structures in experiments. The Si-III structure has been reported as a metastable phase, obtained through slow pressure release from the high-pressure $\beta$-Sn phase [76]. The Si-XII structure was also experimentally obtained by pressure release from the $\beta$-Sn phase and exists over a relatively wide pressure range of 2–12 GPa [77]. The structure with space group $Cmcm$ was synthesized using a high-pressure precursor method and $\rm{Na_{4}}\rm{Si_{24}}$ precursor [78]. The Si-IV structure [79] was synthesized by heating the Si-III structure at 200–600 C [80].
Figure 10 presents the relative enthalpy values for local minimum structures computed by DFT calculations. Structures with relative enthalpy values exceeding 10 meV/atom over the entire pressure range are excluded from Fig. 10. In addition to the stable structures listed in Table 1, three metastable local minimum structures are identified in the 0–80 GPa range in Fig. 10(a). They correspond to the Si-IV structure, a structure with space group $I4_{1}/acd$ represented with 32 atoms, and a structure with space group $R\bar{3}m$. Moreover, it is also observed that the enthalpy values of the Si-XI structure are nearly identical to those of the $\beta$-Sn-type structures. Between 80 and 100 GPa in Fig. 10(b), 15 metastable structures with low relative enthalpy values are found. These structures exhibit different stacking orders of close-packed planes, indicating that elemental Si tends to adopt close-packed structures under high pressures.
@@ IV Phase stability at finite temperature
In this section, we investigate the phase stability at finite temperatures for the local minimum structures identified through the global structure search in elemental Si. To evaluate their free energies, we employ the SSCHA method. The SSCHA method provides a more accurate evaluation of free energy at high temperatures compared to the harmonic approximation. Moreover, this method can predict the dynamical stability of structures that are dynamically unstable under the harmonic approximation.
@@ IV.1 SSCHA formulation
where $F_{\bm{\theta}}$ represents the harmonic expression of the free energy, which is calculated using the effective harmonic force constants (FCs) ${\bm{\theta}}$. $V_{\text{BO}}^{(s)}$ and $\mathcal{V}^{(s)}$ correspond to the Born–Oppenheimer energy and the harmonic potential of the $s$-th sample structure, respectively. The differences between these quantities are averaged over $N_{\rm{samp}}$ sample structures, which are generated from the density matrix at temperature $T$ with FCs ${\bm{\theta}}$. A detailed formulation of Eq. (13) and the density matrix can be found in Ref. [45]. The SSCHA effectively captures the temperature dependence of phonon frequencies and incorporates anharmonic contributions into the free energy calculation.
@@
$\bm{\theta}_{{\rm{mix}}{[l]}}=\alpha\bm{\theta}_{[l]}+(1-\alpha)\bm{\theta}_{[l-1]},$ (15)
where the mixing parameter $\alpha$ is set to 0.5 in this study. The current procedure for SSCHA calculations, coupled with polynomial MLPs, is implemented in the pypolymlp code [53, 66]. Calculations based on the FCs, similar to those used in the harmonic approximation, are performed using the phonopy code [67].
@@ IV.2 Computational procedures
Subsequently, the volume dependence of the Helmholtz free energy is computed for the selected structures using the SSCHA over a range of temperatures. For each structure, 10 to 20 distinct volumes are considered, and equilibrium configurations at these volumes are determined using the MLP. In cases where the structure transforms into a different type during MLP-based geometry optimization, the configuration for a given volume is obtained by isotropically adjusting the volume of the structure at a reference volume where it remains stable. SSCHA calculations are conducted using supercells containing between 64 and 216 atoms, and the number of sample structures in each SSCHA iteration is varied between 5,000 and 20,000, depending on the number of FC components. The Helmholtz free energy values obtained from the systematic SSCHA calculations are then fitted to the Rose–Vinet EOS [70]. The temperature range extends from 0 to 1000 K in 50 K intervals. In total, Helmholtz free energy values are computed for 23,583 distinct combinations of structures, volumes, and temperatures. Notably, the polynomial MLP significantly accelerates the energy and force calculations for large sample sets and across a wide range of conditions.
@@
In these SSCHA calculations, we initially employ the optimal MLP used in the global structure search, which exhibits reasonable accuracy for SSCHA calculations. To further improve accuracy, we then develop an MLP specifically optimized for precise predictions of the target structure. This MLP is henceforth referred to as the specialized MLP, while the MLP used in the global structure search is referred to as the general-purpose MLP to distinguish between the two. Due to the high computational cost associated with the DFT calculations required for training, the development of specialized MLPs is limited to 26 structures. These structures are selected based on their Gibbs free energy values, which are within 5 meV/atom of the lowest Gibbs free energy among the set of 81 local minimum structures.
@@
For each structure, we construct a DFT dataset comprising 150–300 randomly generated sample structures using the effective harmonic FCs. Only those FCs that have converged through SSCHA calculations with an MLP and that yield dynamically stable states are utilized. These sample structures are selected to ensure that the specialized MLP adequately covers the entire ranges of volume and temperature. DFT calculations for the sample structures were carried out using the vasp code [63, 64, 65] with the same computational settings as those used for constructing the DFT dataset, as detailed in Sec. II.3. The cutoff energy was set to 400 eV. The regression coefficients of the polynomial MLP were determined using weighted linear ridge regression, and leave-one-out cross-validation (CV) was utilized to estimate the prediction error, as described in Appendix A. We constructed 285 polynomial MLPs with different single polynomial MLP models. The specialized polynomial MLP with the lowest CV score was selected for the SSCHA calculations. Using the specialized MLPs, SSCHA calculations were performed over the temperature range from 0 to 1000 K.
@@ IV.3 Phase stability evaluated by the SSCHA
Next, we conduct SSCHA calculations using specialized MLPs to obtain more accurate Gibbs free energy values for 26 local minimum structures. Figure 12 shows the pressure-temperature phase diagram for elemental Si calculated using the SSCHA with the specialized MLPs. The transition pressures at room temperature calculated using the MLPs are similar to the experimental values. At higher temperatures, the pressure ranges where the Si-XI, Si-VII, and $\alpha$-La-type structures are stable expand. The stable structures identified in the phase diagram correspond to those computed at zero temperature, as listed in Table 1.
Figure 13: Relationship between pressure and the number of local minimum structures in elemental Si. Only structures with relative Gibbs free energies below 5 meV/atom are included, excluding the global minimum structure. The pressure ranges from 0 to 100 GPa, and the numbers are shown for temperatures of 200, 500, and 1000 K.
Figure 13 illustrates the relationship between pressure and the number of local minimum structures with relative Gibbs free energies below 5 meV/atom, excluding the global minimum structure. The number of these local minimum structures remains nearly constant as the temperature increases. Additionally, many structures exhibit low relative Gibbs free energies in the 80–100 GPa range, where close-packed structures such as the HCP-, $\alpha$-La-, and FCC-type structures are stable. The maximum number of local minimum structures with relative Gibbs free energies below 5 meV/atom is 17, observed at 86.5 GPa and 900 K. Among these structures, 16 correspond to those with low relative enthalpy, as shown in Fig. 10, in the pressure range of 80–100 GPa at 0 K.

### v2

@@ III.1 Methodology and search space
We employ an RSS method to enumerate globally stable and metastable structures, which corresponds to the multi-start method in global optimization [79]. In this approach, local geometry optimizations are repeatedly performed on random initial structures, and the local minimum structure with the lowest enthalpy is considered the global minimum. This heuristic approach has been previously utilized in the ab initio random structure search (AIRSS) [80, 81, 2, 82].
@@ III.2 Iterative update of MLP
We utilize the polynomial MLP selected in Sec. II.5 for the global structure search. This MLP is selected from among Pareto-optimal MLPs to balance prediction accuracy for energy and computational efficiency. Although the MLP exhibits high predictive power for various structures, it occasionally predicts incorrect local minimum structures and fails to accurately estimate their energy values.
To enhance the reliability of the global structure search obtained from the MLP, the MLP is iteratively updated through repeated global structure searches, as follows. Single-point DFT calculations are performed on all local minimum structures identified by the RSS with the MLP, and these results are added to the training dataset. We then retrain the MLP with this expanded dataset and repeat the global structure search using the updated MLP. This iterative approach, which combines the polynomial MLP and RSS, has been previously employed in global structure searches utilizing the polynomial MLP [36, 37, 38].
@@ III.4 Accuracy of updated MLP
Figure 4: (a) Distributions of $E+\bar{\sigma}V$ for local minimum structures, computed using both the DFT calculation and the polynomial MLP, at the first iteration of the iterative MLP update procedure. (b) Distributions of $E+\bar{\sigma}V$ for local minimum structures, computed using the DFT calculation and the polynomial MLP, at the final iteration. For comparison, distributions computed using various interatomic potentials are also shown, including the MEAM potential [43], the Tersoff potential [41], the quadratic SNAP [44], and the GAP [18]. The local minimum structures used in these evaluations are obtained from the RSS with the polynomial MLP at the final iteration. In the figure, “Poly. MLP” and “Quad. SNAP” denote the polynomial MLP and the quadratic SNAP, respectively.
@@
Figure 4(a) and 4(b) show the distributions of $E+\bar{\sigma}V$ for local minimum structures obtained in the first and final iterations, evaluated by the polynomial MLPs and DFT calculations. The distribution for each iteration is obtained using the polynomial MLP applied at the corresponding iteration. Here, $E$ represents the cohesive energy, $V$ denotes the cell volume, and $\bar{\sigma}$ is the mean normal stress, defined as $\bar{\sigma}=(\sigma_{11}+\sigma_{22}+\sigma_{33})/3$ [83]. In the first iteration, many local minimum structures exhibit large prediction errors. However, by incorporating these structures and their DFT-computed properties into the training dataset and retraining the polynomial MLP, the predictive accuracy for local minimum structures improves significantly in the final iteration.
@@
The enthalpy of a local minimum structure is given by $E+pV$ when the stress tensor satisfies $\sigma_{\alpha\beta}=p\delta_{\alpha\beta}$, where $\delta_{\alpha\beta}$ is the Kronecker delta. In Fig. 4, the stress tensor of the local minimum structure optimized by the polynomial MLP is evaluated through a single-point DFT calculation. Because of prediction errors in the polynomial MLP, DFT stress tensor components deviate from the condition $\sigma_{\alpha\beta}=p\delta_{\alpha\beta}$. Consequently, the DFT enthalpy value of the local minimum structure cannot be directly obtained from a single-point DFT calculation. Therefore, Fig. 4 instead presents the distributions of the alternative measure $E+\bar{\sigma}V$, which serves as an approximate evaluation of the predictive accuracy for enthalpy.
III.5 Comparison with existing potentials
Figure 4(b) also exhibits the distributions of $E+\bar{\sigma}V$ for local minimum structures obtained using the polynomial MLP in the final iteration, as evaluated with various interatomic potentials. These include the modified embedded atom method (MEAM) potential [43], the Tersoff potential [41], the quadratic spectral neighbor analysis potential (SNAP) [44], and the Gaussian approximation potential (GAP) [18]. All of these potentials fail to accurately predict the energies of a substantial number of the local minimum structures. The mean absolute errors (MAEs) in $E+\bar{\sigma}V$ exceed 5000 meV/atom for the MEAM potential, the Tersoff potential, and the quadratic SNAP. In contrast, the GAP yields a lower MAE of 267 meV/atom, while the polynomial MLP achieves a substantially reduced MAE of only 19 meV/atom. Moreover, a rough estimate of computational efficiency suggests that the polynomial MLP is approximately three times more efficient than the GAP. Such an MLP with low prediction errors for local minimum structures is essential for global structure searches.
Note that, in these comparisons, the properties obtained from our DFT calculations are used as reference values. While the current polynomial MLP was trained on datasets generated from our own DFT calculations, the other interatomic potentials were developed using different reference datasets based on their respective methodologies. The predictions of interatomic potentials derived from DFT data depend on the computational methods and settings employed, such as the choice of exchange–correlation functional. Moreover, when potentials are constructed based on experimental data, their predictions may deviate from those obtained within our DFT framework. Consequently, it is challenging to make entirely fair comparisons between the predictive performance of our polynomial MLP and that of other potentials. Nevertheless, since our DFT calculations follow widely accepted computational settings, it is reasonable to conclude that our polynomial MLP predicts the properties of a broad range of local minimum structures with an accuracy comparable to that of conventional DFT calculations.
III.6 Globally stable and metastable structures
@@
By performing the RSS with the updated MLP for elemental Si, we enumerate a total of 28,372 local minimum structures. In the final iteration, the total number of energy and force evaluations required for geometry optimizations of random structures using the MLP is 3,062,335,545. The MLP achieves accurate and highly efficient evaluations that would be computationally prohibitive with the DFT. Figure 5(a) presents their relative enthalpy values computed by the MLP. At each pressure, the relative enthalpy is defined with respect to the lowest enthalpy value among the local minimum structures found at the corresponding pressure. The MLP enables the enumeration of numerous structures across all pressure settings. However, the enthalpy values predicted by the MLP still exhibit small but non-negligible prediction errors. Therefore, we compute their enthalpy values using DFT calculations to evaluate the phase stability among these local minimum structures, as follows. First, we select 344 local minimum structures with MLP-calculated relative enthalpies below 30 meV/atom, as shown in Fig. 5(a). We then perform DFT-based geometry optimizations for these structures and eliminate duplicates.
@@
Table 2 lists the metastable local minimum structures, which correspond to experimentally reported structures [90, 91, 92, 93]. Although the Si-XI structure is not identified in the current global structure search, as mentioned above, all other experimentally reported structures are discovered. We evaluate their enthalpy values using DFT calculations if the MLP predicts enthalpies below 30 meV/atom; otherwise, we use the MLP-calculated values. These structures exhibit positive relative enthalpy values, which is consistent with the fact that they have not been observed as stable structures in experiments. The Si-III structure has been reported as a metastable phase, obtained through slow pressure release from the high-pressure $\beta$-Sn phase [90]. The Si-XII structure was also experimentally obtained by pressure release from the $\beta$-Sn phase and exists over a relatively wide pressure range of 2–12 GPa [91]. The structure with space group $Cmcm$ was synthesized using a high-pressure precursor method and $\rm{Na_{4}}\rm{Si_{24}}$ precursor [92]. The Si-IV structure [93] was synthesized by heating the Si-III structure at 200–600 C [94]. Other metastable structures with low relative enthalpy values are shown in the supplemental material, and many energetically metastable structures appear in the pressure range of 80 to 100 GPa.
@@ IV Phase stability at finite temperature
In this section, we present a systematic procedure for performing SSCHA calculations on local minimum structures identified through global structure search. This procedure is applied to assess the phase stability of elemental silicon at finite temperatures. The SSCHA method enables a more accurate estimation of the free energy at elevated temperatures compared to the harmonic approximation. Furthermore, it allows for the prediction of dynamical stability in structures that appear dynamically unstable within the harmonic framework.
The SSCHA can be straightforwardly applied to a single structure with many symmetry operations, as demonstrated in the literature [54, 95, 96]. However, the systematic application of the SSCHA to a wide range of local minimum structures remains challenging for several reasons. (1) A set of local minimum structures usually includes many configurations with lower symmetry. In such cases, SSCHA calculations require the optimization of numerous independent force constant (FC) components. A projector-based approach, recently developed by one of the authors, can significantly accelerate the estimation of these components [97], thereby enabling efficient optimization. (2) SSCHA calculations under various temperature and volume conditions are essential for evaluating the pressure- and temperature-dependent free energy of each local minimum structure. (3) Since SSCHA calculations must be performed for many local minimum structures, the total number of required computations becomes extremely large. (4) To efficiently manage this computational demand, it is essential to employ MLPs that accelerate energy and force evaluations during SSCHA iterations. To ensure sufficient accuracy in determining relative phase stability, MLPs must be developed that can accurately represent both the local minimum structures and configurations with large atomic displacements. The application of the SSCHA in this study establishes a robust procedure for systematic SSCHA calculations supported by MLPs.
@@ IV.1 SSCHA formulation
where $F_{\bm{\theta}}$ represents the harmonic expression of the free energy, which is calculated using the effective harmonic FCs ${\bm{\theta}}$. $V_{\text{BO}}^{(s)}$ and $\mathcal{V}^{(s)}$ correspond to the Born–Oppenheimer energy and the harmonic potential of the $s$-th sample structure, respectively. The differences between these quantities are averaged over $N_{\rm{samp}}$ sample structures, which are generated from the density matrix at temperature $T$ with FCs ${\bm{\theta}}$. A detailed formulation of Eq. (14) and the density matrix can be found in Ref. [54]. The SSCHA effectively captures the temperature dependence of phonon frequencies and incorporates anharmonic contributions into the free energy calculation.
@@
$\bm{\theta}_{{\rm{mix}}{[l]}}=\eta\bm{\theta}_{[l]}+(1-\eta)\bm{\theta}_{[l-1]},$ (16)
where the mixing parameter $\eta$ is set to 0.5 in this study. The current procedure for SSCHA calculations, coupled with polynomial MLPs, is implemented in the pypolymlp code [65, 78]. Calculations based on the FCs, similar to those used in the harmonic approximation, are performed using the phonopy code [99].
@@ IV.2 Computational procedures
Subsequently, the volume dependence of the Helmholtz free energy is computed for the selected structures using the SSCHA over a range of temperatures. For each structure, 10 to 20 distinct volumes are considered, and equilibrium configurations at these volumes are determined using the MLP. In cases where the structure transforms into a different type during MLP-based geometry optimization, the configuration for a given volume is obtained by isotropically adjusting the volume of the structure at a reference volume where it remains stable. SSCHA calculations are conducted using supercells containing between 64 and 216 atoms, and the number of sample structures in each SSCHA iteration is varied between 5,000 and 20,000, depending on the number of FC components. The Helmholtz free energy values obtained from the systematic SSCHA calculations are then fitted to the Rose–Vinet EOS [84]. The temperature range extends from 0 to 1000 K in 50 K intervals.
As shown in Fig. 6, systematic SSCHA free energy calculations are performed twice within the workflow. In the first stage, we utilize the polynomial MLP selected in Sec. II.5, which was also employed in the global structure search and demonstrates reasonable accuracy for SSCHA calculations. In total, Helmholtz free energies are evaluated for 23,583 unique combinations of structures, volumes, and temperatures. The total number of energy and force evaluations required for SSCHA calculations, performed on supercells containing between 64 and 216 atoms, amounts to 1,370,150,000. The use of the polynomial MLP drastically reduces the computational cost of these evaluations, thereby enabling SSCHA calculations over a broad range of local minimum structures that would be computationally prohibitive with DFT.
@@
In the second stage, to further improve accuracy, we then develop an MLP specifically optimized for precise predictions of the target structure, following a similar approach to that used in Ref. [100] for the development of on-the-fly MLPs. This MLP is henceforth referred to as the specialized MLP, while the MLP used in the global structure search is referred to as the general-purpose MLP to distinguish between the two. These specialized MLPs are not suitable replacements for general-purpose MLPs, except when used for SSCHA calculations, because their accuracy decreases for structures other than the target one. However, they allow for more reliable free-energy calculations than general-purpose MLPs. While developing these specialized MLPs requires additional computational resources, their better performance in the large number of energy and force calculations needed for SSCHA justifies the extra computational cost. To limit the additional computational cost associated with the DFT calculations required for training, the development of specialized MLPs is restricted to 26 structures. These structures are selected based on their Gibbs free energy values, which are within 5 meV/atom of the lowest Gibbs free energy among the set of 81 local minimum structures.
@@
For each structure, we construct a DFT dataset comprising 150–300 randomly generated sample structures using the effective harmonic FCs. Only those FCs that have converged through SSCHA calculations with an MLP and that yield dynamically stable states are utilized. These sample structures are selected to ensure that the specialized MLP adequately covers the entire ranges of volume and temperature. DFT calculations for the sample structures were carried out using the vasp code [75, 76, 77] with the same computational settings as those used for constructing the DFT dataset, as detailed in Sec. II.3. The cutoff energy was set to 400 eV. The regression coefficients of the polynomial MLP were determined using weighted linear ridge regression, and leave-one-out cross-validation (CV) was utilized to estimate the prediction error, as described in Appendix C. We constructed 285 polynomial MLPs with different single polynomial MLP models. The specialized polynomial MLP with the lowest CV score was selected for the SSCHA calculations. Using the specialized MLPs, SSCHA calculations were performed over the temperature range from 0 to 1000 K.
The SSCHA calculations with the specialized MLPs require a total of 469,230,000 energy and force evaluations. In the entire SSCHA workflow, 1,839,380,000 supercell energy and force evaluations are performed using MLPs, whereas only 5,910 DFT energy and force evaluations are needed to develop the specialized MLPs. This on-the-fly training strategy enables accurate SSCHA calculations while significantly reducing the computational cost of DFT calculations.
@@ IV.3 Phase stability evaluated by the SSCHA
Next, we conduct SSCHA calculations using specialized MLPs to obtain more accurate Gibbs free energy values for 26 local minimum structures. Figure 7 shows the pressure-temperature phase diagram for elemental Si calculated using the SSCHA with the specialized MLPs. The transition pressures at room temperature calculated using the MLPs are similar to the experimental values. At higher temperatures, the pressure ranges where the Si-XI, Si-VII, and $\alpha$-La-type structures are stable expand. The stable structures identified in the phase diagram correspond to those computed at zero temperature, as listed in Table 1. The number of local minimum structures with low relative Gibbs free energies is provided in the supplemental material, and this number remains nearly constant with increasing temperature. Many structures with low relative Gibbs free energies are found in the 80 to 100 GPa range.